OpenStudy (anonymous):

Is this right? Check my answer please? Question and answer i think is right is in comment.

OpenStudy (anonymous):

\[\frac{10a^2}{ 7b^2}\div \frac{25a^3}{14b^3} \]

OpenStudy (anonymous):

I think the answer is...

OpenStudy (anonymous):

\[\frac{ 5a }{ 4b }\]

OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

@Mertsj ?

OpenStudy (e.mccormick):

gimmi a min. hehe

OpenStudy (anonymous):

ok.

OpenStudy (e.mccormick):

Did you flip the first one?

OpenStudy (anonymous):

let me re-check.

OpenStudy (e.mccormick):

\[\frac{a}{b}\div \frac{c}{d}=\frac{a}{b}\times \frac{d}{c}\]But what I see your answer as being would be if you flipped the a/b and not the c/d.

OpenStudy (anonymous):

one sec handling sumthin over b right with u.

OpenStudy (anonymous):

ok, had a call

OpenStudy (anonymous):

What do you mean?

OpenStudy (e.mccormick):

Well, what did you do to divide?

OpenStudy (anonymous):

I divided 7 by 14 and got 2

OpenStudy (anonymous):

Will it have a ^5?

OpenStudy (e.mccormick):

AH! Does not work that way. Dividing one fraction by another is the same as multiplying the first fraction by the reciprocal of the second. AKA: Flip the second and multiply.

OpenStudy (anonymous):

Can you draw that process please?

OpenStudy (anonymous):